OpenStudy (anonymous):

Check whether (– 1, 5) is a solution of each system of equations. 2x=y-7 2y=-x+9

OpenStudy (supernova_sonntag):

x=-1 y=5 Plug in your values: 2(-1)=5-7 Does -2= -2? Yes. Solution. 2(5)=(1)+9 {Because x is already negative, I made it positive in the problem.} Does 10=10? Yes. Solution.